DTERMINE IF (1/3,-5)is a solution-3x-10y=51
Answer by Theo(13342)

your equation is -3x - 10y = 51
x = (1/3)
y = (-5)
substitute in the equation to get:
-3 * (1/3) - (10 * -5) = 51
simplify to get:
-3/3 - (-50) = 51
simplify further to get:
-1 + 50 = 51
combine like terms to get:
49 = 51
since this is not true, then those values are not solutions to the equation.
if your equation was:
3x - 10y = 51, then:
substituting would get you:
3*(1/3) - (10*(-5)) = 51 which would simplify to:
1 - (-50) = 51 which would simplify further to:
1 + 50 = 51 which would become:
51 = 51
in that case it would be a solution to the equation.
it depends on what your equation is.
the process is the same.
the result is different if i assume your original equation is:
-3x-10y=51 or whether i assume your original equation is:
3x-10y = 51
check your original equation to make sure what it is and you can find the answer depending on that.
the process is the same regardless of the equation.
you replace the variable x with its value and the replace the variable y with its value and then you solve.
the coordinate pair of (1/3, -5) is the (x,y) coordinate pair which means that the value of x is equal to 1/3 and the value of y is -5.